225 al 93% 12:22 PM webwork.phys.yorku.ca/webwc THE PROBLEM A bar of length L 3.6 m is balanced at is midpoint. Then a person of mass m-29 kg sits on one end. Where should a person of mass M-63 kg sit so that the bar will once again be balanced at its midpoint? PAPER SOLUTION INTERPRET Identify all of the true statements A. The center of mass of the bar itself is always at the center of the bar throughout this question. B. The center of mass of the total system is only at the bars center when the bar is balanced. c. The bar is balanced if and only if the center of mass of the pair of people is at the center of the bar □ D. None of the above. DEVELOP Derive an algebraic expression for the distance between the people when the bar is balanced.
Answer #1

COnsider Moment about the pivot,
\sum M_{0} = 0

29*(\frac{3.6}{2})-69*x = 0

Where x is the distance of the second person from the midpoint pivot

x = 0.7565 m

Distance between the two persons,
d = \frac{L}{2}+x

d = \frac{3.6}{2}+0.7565

d = 2.5565 m